you have $50,000 to invest, you invest some at 5% and the rest at 4%. you earn $2100 in annual interest, how much did you invest at each rate?
Answer by Flake(45)

Let X = money invest at 5%
Let Y = money invest at 4%
-------Step#1
X+Y = 50,000
Y = 50,000 - X
-------Step#2
(5%)X + (4%)Y = $2100
0.05X + 0.04(50,000-X) = 2100
0.01X = 100
X = $10,000
Hence Y = $40,000